OpenStudy (anonymous):

I need help on how to solve: the regular price of a tv is $200 but it's on sale for 25% off what is the sale price?

OpenStudy (anonymous):

That is the discount. So, sale price is $150.

OpenStudy (anonymous):

so its 150

OpenStudy (anonymous):

That is what I got but it doesn't show that on the riddle... ( it's on a work sheet)

OpenStudy (anonymous):

Oh wait never mind but how did u get 150

OpenStudy (anonymous):

Sale price = Original Price - Discount Value Sale price = 200 - 200 * 25% Sale price = 200 - 200 * 0.25 Sale price = 200 - 200 * 1/4 Sale price = 200 - 200/4 Sale price = 200 - 50 Sale price = 150

OpenStudy (anonymous):

Oooooooh thanks! 😋